OpenStudy (anonymous):

What is the solution for x^2 + 4x = 96?

OpenStudy (anonymous):

Are the solutions***

OpenStudy (campbell_st):

rewrite the equation as \[x^2 + 4x - 96 = 0\] the quadratic can be solved by factoring... find the factors of -96 that add to 4 the larger factor is positive... then its (x + factor 1)(x + factor 2) = 0 hope it helps

OpenStudy (anonymous):

The factors are 12 and 8?

OpenStudy (anonymous):

Add to for though...

OpenStudy (anonymous):

it***

OpenStudy (campbell_st):

nearly... so its 12 and -8 so you have (x +12)(x -8) = 0 so if either binomial is zero, the equation is zero so you need to solve for x, the following x + 12 = 0 and x - 8 = 0 and then you'll have your solutions.

OpenStudy (anonymous):

-12 and 8?

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

thats correct